OpenStudy (anonymous):

0.3x + 0.2y = 5, 0.5x + 0.4y =11

OpenStudy (anonymous):

Eqn 1 0.3x + 0.2y = 5 0.2y = 5 - 0.3x 0.4y = 10 - 0.6x Eqn 2 0.5x + 0.4y = 11 0.4y = 11 - 0.5x 10 - 0.6x = 11 - 0.5x Substitute Eqn in to eqn 2 10 - 11 = -0.5x + 0.6x rearrange -1 = 0.1x -10 = x x = -10 0.4y = 11 - 0.5(-10) = 11 + 5 = 16 0.4y = 16 (10/4)(4/10)y = y = (10/4)(16) = 40 checking: x = -10, y = 40 0.5(-10) + 0.4(40) = -5 + 16 = 11 Correct!